### v1

# Conserved quantities of distinguished curves on conformal sphere

Abstract.
We give conserved quantities of two generalizations of conformal circles on the conformal sphere. One generalization concerns curves carrying a distinguished parallel tractor along them, which can be used to construct the conserved quantities. The other is the class of curves satisfying the conformal Mercator equation, the conserved quantities of which are computed using Lagrangian formalism. The two generalizations are not disjoint, and our main result is the relation between their conserved quantities. Explicit examples are also presented.
Key words and phrases:
conserved quantities; distinguished curves; conformal tractors; conformal Mercator equation; conformal round sphere;
1. Introduction
By distinguished curves on a conformal manifold, we mean a class of curves that is invariant under conformal transformations. A well-known example of these is the class of conformal circles (or conformal geodesics), whose definition could be traced back to Yano [16] who defined them as solutions of a third order ODE. Other equivalent definitions in terms of the defining ODEs are given in [2, 15]. There are also extensive studies of conformal circles based on tractor formulation, which can be found, for example, in [3, 14, 8] and references therein.
This paper focuses on distinguished curves on the conformal sphere, which is the round $n$-sphere with the conformal class of the standard round metric. The conformal sphere is a fundamental example of conformal manifolds. In particular, it admits maximum conformal symmetries and is a homogeneous manifold. Conformal circles on the conformal sphere are very well known, in fact they are all round circles. Another example of distinguished classes of curves is the family of conformal loxodromes, recently defined in [7]. Our work here concentrates on two generalizations of conformal circles, and the computation of their conserved quantities. Recall that the conserved quantities for conformal circles have previously been studied, for example, in [15, 14, 6, 9]. In [15, 6], conserved quantities were constructed from conformal Killing-Yano $2$-forms and then used to prove complete integrability of the conformal circle equation on gravitational instantons. On the other hand, in [14] tractor method was used to construct conserved quantities from conformal Killing vector fields.
Our first generalization follows [9], where authors use tractor calculus to construct conserved quantities of conformal circles. Explicitly, they construct a series of canonical tractors defined for each curve and show that the wedge of the first three of them is parallel along the curve only if the curve is a conformal circle. Then pairing with suitable BGG solutions gives conserved quantities of conformal circles. We focus in Section 2 on curves such that the wedge of the first four canonical tractors is parallel. On the conformal round sphere, it is then a direct generalization to find conserved quantities for such curves by pairing with suitable parallel tractors.
The second generalization is based on a Lagrangian formulation for the conformal circle equation. It was shown in [5] that the equation, on an arbitrary conformal manifold, can be derived from a conformally invariant $3$rd order Lagrangian, provided that one considers a special class of variations. [^1] Variational formulation of conformal circle equation has also recently been studied in [12, 11], where it is shown that unparametrised conformal circles in $3$ dimensions is precisely the Euler-Lagrange equation of the total torsion functional. The Euler-Lagrange equation of this Lagrangian, on the other hand, is of $4$th order, and called the conformal Mercator equation in [5]. In fact, the equation also arises from the $2$nd order Lagrangian discussed in [2], which differs from the $3$rd order Lagrangian of [5] by a total derivative term. It was noted [2, 5] that on a conformally-flat manifold such as the conformal sphere, its solutions cover conformal circles. Thus, we shall consider parametrized curves which are solutions of the conformal Mercator equation as another class of distinguished curves generalizing the conformal circles. [^2] Relation between solutions of the conformal Mercator equation and conformal loxodromes, as defined in [7], has been noted in [7]. As the associated Lagrangian is conformally invariant, we are able to use Noether’s principle to derive conserved quantities for the solution curves from conformal Killing vector fields. This result is given in Section 3.
Via an explicit example, it is evident that the two generalizations we consider are not disjoint, that is, they share some common curves on which the two sets of our conserved quantites are both constant. This motivates us to find a relation between the generally different quantities. Using the Hamiltonian formulation which simplifies the expressions of the quantities, we establish such a relation. This is our main result and is given in Section 4.
2. Invariants via parallel tractors
We use here tractor calculus to study parallel tractors and conserved quantities of curves on the conformal sphere. One can see e.g. [4] for an introduction to conformal tractors.
2.1. Conformal sphere and tractors
By a conformal sphere we mean the sphere $S^{n}$ with the conformal class of metrics represented by the standard round metric. It is a homogeneous space and in fact, maximally symmetric conformal Riemannian structure of dimension $n$. The Lie group of conformal transformations of $S^{n}$ is isomorphic to $O(n+1,1),$ with the corresponding Lie algebra $\mathfrak{so}(n+1,1).$ This follows from the realization of $S^{n}$ as the projectivization of the cone of non-zero null-vectors in the pseudo-Euclidean space $\mathbb{R}^{n+1,1}$ of signature $(n+1,1)$. The group $O(n+1,1)$ acts transitively on $S^{n}$ and the stabilizer of a point is the Poincaré subgroup $P\subset O(n+1,1)$, so $S^{n}\cong O(n+1,1)/P$. We write elements of $\mathfrak{so}(n+1,1)$ as $(1,n,1)$-block matrices
$\left(\begin{smallmatrix}-a&2S&0\\ T&-R&-2S^{t}\\ 0&-T^{t}&a\end{smallmatrix}\right)$
for $a\in\mathbb{R}$, $R\in\mathfrak{so}(n)$ and $T,S^{t}\in\mathbb{R}^{n}$. Viewed as a conformal Killing field, we get
$(T^{i}-{R^{i}}_{j}x^{j}+ax^{i}-2S_{j}x^{j}x^{i}+x_{j}x^{j}S^{i})\frac{\partial}{\partial x^{i}}.$
The Lie algebra $\mathfrak{p}\subset\mathfrak{so}(n+1,1)$ corresponds to the upper triangular matrices and there is a natural complement $\mathfrak{c}$ of $\mathfrak{p}$ in $\mathfrak{so}(n+1,1)$ such that $\mathfrak{so}(n+1,1)=\mathfrak{c}\oplus\mathfrak{p}$, i.e.
$\mathfrak{p}\ni\left(\begin{smallmatrix}-a&2S&0\\ 0&-R&-2S^{t}\\ 0&0&a\end{smallmatrix}\right),\ \ \mathfrak{c}\ni\left(\begin{smallmatrix}0&0&0\\ T&0&0\\ 0&-T^{t}&0\end{smallmatrix}\right).$
The complement gives natural local exponential coordinates on $S^{n}$ around the origin $o=eP$, which correspond to the usual coordinates coming from the projectivization of the cone in $\mathbb{R}^{n+1,1}$. Indeed, we get
$\exp\left(\begin{smallmatrix}0&0&0\\ T&0&0\\ 0&-T^{t}&0\end{smallmatrix}\right)\left(\begin{smallmatrix}1\\ 0\\ 0\end{smallmatrix}\right)=\left(\begin{smallmatrix}1&0&0\\ T&E&0\\ -\frac{1}{2}\langle T,T\rangle&-T^{t}&1\end{smallmatrix}\right)\left(\begin{smallmatrix}1\\ 0\\ 0\end{smallmatrix}\right)=\left(\begin{smallmatrix}1\\ T\\ -\frac{1}{2}\langle T,T\rangle\end{smallmatrix}\right),$
[…]
We employ the tractor point of view in Section 2, and we can follow [10] due to the homogeneity of the conformal sphere. We consider the standard tractor bundle $\mathcal{T}$ modeled on $\mathbb{R}^{n+2}$ with the tractor metric, which is the standard pseudometric of signature $(n+1,1)$. Thus, we deal with the standard representation
$\rho:\mathfrak{so}(n+1,1)\to\mathfrak{gl}(\mathbb{R}^{n+2}),$
and we employ the description of the tractors in the above exponential coordinates just by restriction to $\mathfrak{c}$. Indeed, the standard tractor bundle $\mathcal{T}$ trivializes as $\mathcal{T}\simeq\mathfrak{c}\times\mathbb{R}^{n+2}$ and we can see the standard tractors as sections of $\mathfrak{c}\times\mathbb{R}^{n+2}$, that is, functions $\mathfrak{c}\to\mathbb{R}^{n+2}$. Then, written in this local trivialization, the corresponding functions take the form
$\exp(J)\mapsto\exp(-\rho(J))(w){\rm\ \ for\ \ }J=\left(\begin{smallmatrix}0&0&0\\ X&0&0\\ 0&-X^{t}&0\end{smallmatrix}\right)$ (2.1)
for some neighborhood $X$ of 00 in $\mathfrak{c}$ and $w=(w^{a})$ are coordinates in the representation space $\mathbb{R}^{n+2}$. Viewing $\rho(J)$ as a linear transformation, we get
$e_{0}\mapsto\sum_{i=1}^{n}X^{i}e_{i},\ \ e_{i}\mapsto-X^{i}e_{N},\ \ e_{N}\mapsto 0,$ (2.2)
where $X=(X^{i})$, $i=1,\dots,n$ and $N:=n+1$. Thus in coordinates $w$ we get
$\rho(J)(w)=w^{0}\sum_{i=1}^{n}X^{i}e_{i}+\sum_{i=1}^{n}w^{i}X^{i}e_{N}.$
One can then compute directly
$w-\rho(J)(w)+\frac{1}{2}\rho^{2}(J)(w),$ (2.3)
and the series stops as $\rho^{3}(J)$ vanishes. The same ideas apply analogously to arbitrary tractor bundles just by tensoriality; we are particularly interested in $\wedge^{4}\mathbb{R}^{n+2}$.
Let us finally say that the standard tractor connection decomposes into the fundamental derivative $D$ and the algebraic action and in the case of the conformal sphere takes the form $\nabla^{\mathcal{T}}=D+\rho$, restricted to $\mathfrak{c}$ to get a coordinate description. We then get the tractor connection on an arbitrary tensor bundle analogously by taking the corresponding representation of $\mathfrak{so}(n+1,1)$.
2.2. Curves and tractors
We consider arbitrary curve $\gamma=(\gamma^{i}(t))$, $i=1,\dots,n$, and we denote
$U=(\frac{d}{dt}\gamma(t)),\ \ A=(\frac{d^{2}}{dt^{2}}\gamma(t)),\ \ A^{\prime}=(\frac{d^{3}}{dt^{3}}\gamma(t)),\dots$
its velocity vector field, acceleration vector field, and so on for subsequent derivatives, where $\,{}^{\prime}$ denotes $\frac{d}{dt}.$ We denote by $u$ the length of the velocity vector field $u=\sqrt{\langle U,U\rangle}$ for the standard scalar product on $\mathbb{R}^{n}$.
For each curve $\gamma$, we define a canonical series of derived standard tractors by means of the tractor derivative. Explicitly, the tractor derivative takes the form $\nabla^{\mathcal{T}}(\mathbb{X})=D(\mathbb{X})+K\cdot\mathbb{X},$ for the tractor $\mathbb{X}$, where
$K:=\left(\begin{smallmatrix}0&0&0\\ U&0&0\\ 0&-U&0\end{smallmatrix}\right).$
We write these tractors in coordinates in the standard basis $e_{a},$ $a=0,\dots,N,$ of $\mathbb{R}^{n+2}$ that we decompose (for the canonical metric from the conformal class) as
$\mathbb{R}^{1+n+1}=\langle e_{0}\rangle+\langle e_{1},\dots,e_{n}\rangle+\langle e_{N}\rangle.$
We start with the canonical inclusion $\mathbb{T}=u^{-1}e_{0}$. This leads to the following sequence of standard tractors, where we use Einstein summation notation and $i=1,\dots,n$
$\mathbb{U}=-u^{-3}\langle U,A\rangle e_{0}+u^{-1}U^{i}e_{i}$
$\mathbb{A}=(3u^{-5}\langle U,A\rangle^{2}-u^{-3}(\langle A,A\rangle+\langle U,A^{\prime}\rangle))e_{0}+(-2u^{-3}\langle U,A\rangle U^{i}+u^{-1}A^{i})e_{i}-ue_{N}$
$\mathbb{A}^{\prime}=(-15u^{-7}\langle U,A\rangle^{3}+9u^{-5}\langle U,A\rangle(\langle A,A\rangle+\langle U,A^{\prime}\rangle)-3u^{-3}\langle A,A^{\prime}\rangle-u^{-3}\langle U,A^{\prime\prime}\rangle)e_{0}$
$+(9u^{-5}\langle U,A\rangle^{2}U^{i}-3u^{-3}(\langle A,A\rangle+\langle U,A^{\prime}\rangle)U^{i}-3u^{-3}\langle U,A\rangle A^{i}+u^{-1}(A^{\prime})^{i})e_{i}$
$\dots$
The best known and studied class of conformal curves are conformal circles, [8, 9, 14]. In the language of tractors, the curve is a conformal circle if and only if tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}}$ are linearly dependent, i.e. $\mathbb{A^{\prime}}=\langle\mathbb{T},\mathbb{U},\mathbb{A}\rangle$. This leads to the condition that $T_{3}=\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ is parallel along $\gamma$ for the tractor connection, [9]. For each (parametrized) curve, thus for each sequence of tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}},\dots$ of arbitrary length $\ell$, we can compute the following Gramm matrix of order $\ell$
$M_{\ell}=\left(\begin{smallmatrix}0&0&-1&0&\alpha_{1}&\cdots\\ 0&1&0&-\alpha_{1}&-\frac{3}{2}\alpha^{\prime}_{1}&\cdots\\ -1&0&\alpha_{1}&\frac{1}{2}\alpha^{\prime}_{1}&\frac{1}{2}\alpha^{\prime\prime}_{1}-\alpha_{2}&\cdots\\ 0&-\alpha_{1}&\frac{1}{2}\alpha^{\prime}_{1}&\alpha_{2}&\frac{1}{2}\alpha^{\prime}_{2}&\cdots\\ \alpha_{1}&-\frac{3}{2}\alpha^{\prime}_{1}&\frac{1}{2}\alpha^{\prime\prime}_{1}-\alpha_{2}&\frac{1}{2}\alpha^{\prime}_{2}&\alpha_{3}&\cdots\\ \vdots&\vdots&\vdots&\vdots&\vdots&\ddots\end{smallmatrix}\right)$
for the standard tractor metric. Then we can use the Cramer rule for $M_{4}$ to write the tractor combination explicitly. Direct computation gives
$\mathbb{A^{\prime}}=-\alpha_{1}\mathbb{U}-\frac{1}{2}\alpha^{\prime}_{1}\mathbb{T},$
where $\alpha_{1}$ is the square of the length of the acceleration tractor $\mathbb{A}$. Then the tractor derivative gives
$\nabla^{\wedge^{3}\mathcal{T}}(\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A})=\mathbb{U}\wedge\mathbb{U}\wedge\mathbb{A}+\mathbb{T}\wedge\mathbb{A}\wedge\mathbb{A}+\mathbb{T}\wedge\mathbb{U}\wedge(-\alpha_{1}\mathbb{U}-\frac{1}{2}\alpha^{\prime}_{1}\mathbb{T})=0.$
This also allows us to define a series of natural relative invariants $\Delta_{\ell}=\det(M_{\ell})$, and it follows that the curve is a conformal circle if and only if $\Delta_{4}=0$. Let us note that $\Delta_{1}=\Delta_{2}=0$ and $\Delta_{3}=-1$, so the case of conformal circles is the first interesting one.
In fact, it is proved in [14, Section 2.3.] that in general, vanishing of $\Delta_{\ell}$ is equivalent to the fact, that the tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}},\dots,\mathbb{A}^{\ell-3}$ are linearly dependent. Let us now generalize to the case of $\Delta_{5}$. Thus we assume $\Delta_{4}\neq 0$ and $\Delta_{5}=0$, which means that the tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A}^{\prime},\mathbb{A}^{\prime\prime}$ are linearly dependent, i.e. $\mathbb{A}^{\prime\prime}\in\langle\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A}^{\prime}\rangle$. Using $M_{5}$ and the Cramer rule, we find the tractor combination explicitly as
$\begin{split}\mathbb{A}^{\prime\prime}&=\frac{\Delta^{\prime}_{4}}{2\Delta_{4}}\mathbb{A}^{\prime}-\alpha_{1}\mathbb{A}+\frac{1}{2\Delta_{4}}(\alpha^{\prime}_{1}(2\alpha_{2}-\Delta_{4})-\alpha_{1}\alpha^{\prime}_{2})\mathbb{U}\\ &+\frac{1}{4\Delta_{4}}(2\alpha_{1}(\alpha^{\prime}_{1})^{2}-4\Delta_{4}^{2}-2\alpha^{\prime\prime}_{2}\Delta_{4}-\alpha^{\prime}_{1}\alpha^{\prime}_{2})\mathbb{T}.\end{split}$ (2.4)
We are interested in analogues of the $3$-tractor $T_{3}$. Let us derive the expression $\Delta_{4}^{a}\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}$ with respect to tractor connection along the curve, where we consider a general power of $\Delta_{4}$ due to the possible effect of the coefficient with $\mathbb{A}^{\prime}$. We get
$\begin{split}&\nabla^{\wedge^{4}\mathcal{T}}(\Delta_{4}^{a}\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}})=a\Delta_{4}^{a-1}\Delta_{4}^{\prime}\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}+\Delta_{4}^{a}\cdot\mathbb{U}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}\\ &+\Delta_{4}^{a}\cdot\mathbb{T}\wedge\mathbb{A}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}+\Delta_{4}^{a}\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A^{\prime}}\wedge\mathbb{A^{\prime}}+\Delta_{4}^{a}\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge(\frac{\Delta^{\prime}_{4}}{2\Delta_{4}}\mathbb{A^{\prime}}+\dots)\\ &=(a\Delta_{4}^{a-1}\Delta^{\prime}_{4}+\Delta_{4}^{a}\cdot\frac{\Delta^{\prime}_{4}}{2\Delta_{4}})\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}.\end{split}$ (2.5)
Thus we get the condition
$a\Delta_{4}^{a-1}\Delta^{\prime}_{4}+\Delta_{4}^{a}\cdot\frac{\Delta^{\prime}_{4}}{2\Delta_{4}}=0$
to get a parallel object. We get $a=-\frac{1}{2}$ or $\Delta^{\prime}_{4}=0$.
Proposition 2.1.
The curve $\gamma(t)$ satisfies $\Delta_{4}\neq 0$ and $\Delta_{5}=0$ if and only if
$(-\Delta_{4})^{-\frac{1}{2}}\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}$ (2.6)
is parallel to the tractor connection along the curve. Moreover, if $\Delta_{4}$ is constant, then already the expression
$T_{4}:=\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}$ (2.7)
is parallel to the tractor connection.
Proof.
The above observations show that the conditions $\Delta_{4}\neq 0$ and $\Delta_{5}=0$ are equivalent to the linear dependency of the tractors $\mathbb{T},\mathbb{U},\mathbb{A},\mathbb{A^{\prime}}$ and $\mathbb{A^{\prime\prime}}$ and the explicit combination takes form (2.4). Note that $\Delta_{4}$ is non-positive, [14, Lemma 2.3.]. The computation (LABEL:derivative-4-tractor) then gives that (2.6) is parallel with respect to the tractor connection. Conversely, assuming (2.6) is parallel, we get
$\nabla^{\wedge^{4}\mathcal{T}}((-\Delta_{4})^{-\frac{1}{2}}\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}})=(-\Delta_{4}^{-\frac{1}{2}})\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge(-\frac{\Delta^{\prime}_{4}}{2\Delta_{4}}\mathbb{A}^{\prime}+\mathbb{A}^{\prime\prime})=0$
and this implies, that $\mathbb{A}^{\prime}$ and $\mathbb{A}^{\prime\prime}$ are dependent modulo $\langle\mathbb{T},\mathbb{U},\mathbb{A}\rangle$. If $\Delta_{4}$ is a non-zero constant, then $\Delta^{\prime}_{4}=0$ and $T_{4}$ is parallel from (LABEL:derivative-4-tractor). ∎
Since $\Delta_{4}$ is a relative invariant, the condition that $\Delta_{4}$ is non-zero constant gives a special class of (parametrized) curves. In particular, we can normalize the constant to $\Delta_{4}=-1$, which corresponds to the conformal arc-length parametrization.
2.3. Conserved quantities and tractors
In [9], authors give a general method, how to use conformal Killing-Yano $2$-forms and the parallel $3$-tractor $\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ to find conserved quantities of conformal circles on arbitrary conformal manifolds. The situation on $S^{n}$ is simpler; let us outline here the situation. For interest, we give the resulting quantities on $S^{n}$ in Appendix A.
On the conformal sphere, each conformal Killing-Yano $2$-form corresponds to an element of the tractor bundle for the representation $\rho_{3}:\mathfrak{so}(n+1,1)\to\mathfrak{gl}(\wedge^{3}\mathbb{R}^{n+2})$. It follows from the tractor theory that each such tractor is parallel with respect to the tractor connection. Indeed, in the case of $S^{n}$, the whole tractor space corresponds to solutions of the corresponding BGG operators, and they are all normal, i.e. coming from sections parallel to the tractor connection. Then pairing each such parallel tractor with the $3$-tractor corresponding to $\gamma$ gives a conserved quantity along the curve $\gamma$. By pairing, we mean the pairing given by the tractor metric on the standard tractor bundle that we represent in the standard basis $e_{0},e_{i},e_{N}$, $i=1,\dots,n$ by the block $(1,n,1)$-matrix
$\left(\begin{smallmatrix}0&0&1\\ 0&E&0\\ 1&0&0\end{smallmatrix}\right)$
for identity matrix $E$. On the conformal sphere, this directly generalizes for curves satisfying $\Delta_{4}\neq 0,\Delta_{5}=0$. Each conformal Killing-Yano $3$-tensor corresponds to an element of the tractor space $\wedge^{4}\mathbb{R}^{n+2}$. In particular, pairing with the $4$-tractor $(-\Delta_{4})^{-\frac{1}{2}}\cdot\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}$ for $\gamma$ gives a conserved quantity along $\gamma$. If, moreover, $\Delta_{4}$ is constant, then pairing with $T_{4}$ already gives a conserved quantity.
Let us describe these quantities explicitly in the basis $e_{abcd}:=e_{a}\wedge e_{b}\wedge e_{c}\wedge e_{d}$ with $0\leq a<b<c<d\leq N$ for the standard basis $e_{a}$, $a=0,\dots,N$. Firstly, we find the $4$-tractor $T_{4}=\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}\wedge\mathbb{A^{\prime}}.$
Proposition 2.2.
The expression $T_{4}$ takes the explicit form
$T_{4}=u^{-4}U^{[i}A^{j}(A^{\prime})^{k]}e_{0ijk}-3u^{-4}\langle U,A\rangle U^{[i}A^{j]}\>e_{0ijN}+u^{-2}U^{[i}(A^{\prime})^{j]}\>e_{0ijN}$ (2.8)
[…]
Proof.
One can see from the explicit forms of tractors $\mathbb{T}$, $\mathbb{U}$, $\mathbb{A}$ and $\mathbb{A^{\prime}}$, that the term with $e_{0}$ comes only from $\mathbb{T}$ and then the remaining tractors can add $e_{i},e_{N}$ only. The tractor $\mathbb{A}$ is the only tractor with $e_{N}$. Moreover, $U\wedge U$ and $A\wedge A$ vanish, so $T_{4}$ is as follows, where we use the Einstein summation notation and $i=1,\dots,n$
$u^{-1}e_{0}\wedge u^{-1}U^{i}e_{i}\wedge\big{(}u^{-1}A^{j}e_{j}-ue_{N}\big{)}\wedge\big{(}-3u^{-3}\langle U,A\rangle A^{k}e_{k}+u^{-1}(A^{\prime})^{k}e_{k}\big{)}=$
$u^{-1}e_{0}\wedge u^{-1}U^{i}e_{i}\wedge u^{-1}A^{j}e_{j}\wedge u^{-1}(A^{\prime})^{k}e_{k}+u^{-1}e_{0}\wedge u^{-1}U^{i}e_{i}\wedge-ue_{N}\wedge-3u^{-3}\langle U,A\rangle A^{k}e_{k}+$
$u^{-1}e_{0}\wedge u^{-1}U^{i}e_{i}\wedge-ue_{N}\wedge u^{-1}(A^{\prime})^{k}e_{k}=$
$u^{-4}e_{0}\wedge U^{i}e_{i}\wedge A^{j}e_{j}\wedge(A^{\prime})^{k}e_{k}+u^{-4}e_{0}\wedge U^{i}e_{i}\wedge e_{N}\wedge 3\langle U,A\rangle A^{k}e_{k}-$
$u^{-2}e_{0}\wedge U^{i}e_{i}\wedge e_{N}\wedge(A^{\prime})^{k}e_{k}.$
The order of wedges does not respect the lexicographic ordering, so we have to reorder them to get a description in the basis $e_{abcd}.$ This leads to the reordering of coefficients, and summarizing them together leads exactly to the anti-symmetrization, so we get the formula for $T_{4}$ in the statement. ∎
Let us now find the conserved quantities. We pair $T_{4}$ with the natural lexicographic basis of the tractor space $\wedge^{4}\mathbb{R}^{n+1}$, where the indices of the quantity reflect the corresponding basis element. In Proposition 2.3 below, recall that the elementary alternating tensor of rank $m,$ $\epsilon^{i_{1}\dots i_{m}},$ is defined by its action on $m$ vectors, $Y,\dots,Z,$ as
$\epsilon^{i_{1}\dots i_{m}}(Y,\dots,Z)=\det\left(\begin{array}[]{ccc}Y^{i_{1}}&\dots&Z^{i_{1}}\\ \vdots&{}\hfil&\vdots\\ Y^{i_{m}}&\dots&Z^{i_{m}}\end{array}\right).$
Proposition 2.3.
The conserved quantities are as follows:
$Q^{0ijN}$ $=3u^{-4}\langle U,A\rangle\epsilon^{ij}(U,A)-u^{-2}\epsilon^{ij}(U,A^{\prime})+u^{-4}\epsilon^{ijl}(U,A,A^{\prime})X_{l}$ (2.9)
$\begin{split}Q^{0ijk}&=3u^{-4}\langle U,A\rangle\epsilon^{ijk}(X,U,A)-u^{-2}\epsilon^{ijk}(X,U,A^{\prime})\\ &+\frac{1}{2}|X|^{2}u^{-4}\epsilon^{ijk}(U,A,A^{\prime})+u^{-4}\epsilon^{ijkl}(X,U,A,A^{\prime})X_{l}\end{split}$ (2.10)
$Q^{ijkN}$ $=u^{-4}\epsilon^{ijk}(U,A,A^{\prime})$ (2.11)
$Q^{ijkp}$ $=u^{-4}\epsilon^{ijkp}(X,U,A,A^{\prime})$ (2.12)
Proof.
The standard action (2.2) differs for $e_{0}$, $e_{i}$ and $e_{N}$. Thus it is reasonable to organize elements of the lexicographic basis of $\wedge^{4}\mathbb{R}^{n+2}$ into the following four types
  1. (1)
    $\frac{1}{2}n(n-1)$ elements of the form $e_{0ijN}$,
  2. (2)
    $\frac{1}{6}n(n-1)(n-2)$ elements of the form $e_{0ijk}$,
  3. (3)
    $\frac{1}{6}n(n-1)(n-2)$ elements of the form $e_{ijkN}$,
  4. (4)
    $\frac{1}{24}n(n-1)(n-2)(n-3)$ elements of the form $e_{ijkl}$,
where $i,j,k,l=1,\dots,n$. For each basis element, i.e. for each coordinate, we compute the expression (2.1) for the action $\rho_{4}:\mathfrak{so}(n+1,1)\to\mathfrak{gl}(\wedge^{4}\mathbb{R}^{n+2})$ given by tensoriality of the standard action. Thus we compute
$w-\rho_{4}(J)(w)+\frac{1}{2}\rho_{4}^{2}(J)(w)$ (2.13)
for coordinates $w$, and we always choose one of the coordinates equal to $1$ and the rest zero. Direct computation using the formula (2.2) and tensoriality give the following actions $\rho_{4}(J)$, where sums are generically for $l=1,\dots,n$ and we emphasize other restrictions below the sum
$e_{0ijN}$ $\mapsto\sum_{l<i}X^{l}e_{lijN}-\sum_{i<l<j}X^{l}e_{iljN}+\sum_{l>j}X^{l}e_{ijlN}$
$e_{0ijk}$ $\mapsto\sum_{l<i}X^{l}e_{lijk}-\sum_{i<l<j}X^{l}e_{iljk}+\sum_{j<l<k}X^{l}e_{ijlk}-\sum_{l<k}X^{l}e_{ijkl}$
$-X^{i}e_{0jkN}+X^{j}e_{0ikN}-X^{k}e_{0ijN}$
$e_{ijkN}$ $\mapsto 0$
$e_{ijkl}$ $\mapsto X^{i}e_{jklN}-X^{j}e_{iklN}+X^{k}e_{ijlN}-X^{l}e_{ijkN}.$
Having these, we compute the second powers aka compositions $\rho_{4}^{2}(J)$ as
$e_{0ijN}$ $\mapsto 0$
$e_{0ijk}$ $\mapsto\sum_{k=1}^{n}\big{(}(X^{l})^{2}-2((X^{i})^{2}+(X^{j})^{2}+(X^{k})^{2})\big{)}e_{ijkN}$
$-2X^{i}(\sum_{l<j,\>l\neq i}X^{l}e_{ljkN}+\sum_{j<l<k}X^{l}e_{jlkN}+\sum_{k<l}X^{l}e_{jklN})$
$+2X^{j}(\sum_{j<i}X^{l}e_{likN}-\sum_{i<l<k,\>l\neq j}X^{l}e_{ilkN}+\sum_{l>k}X^{l}e_{iklN})$
$-2X^{k}(\sum_{l<i}X^{l}e_{lijN}-\sum_{i<l<j}X^{l}e_{iljN}+\sum_{l>j,\>l\neq k}x_{l}e_{ijlN})$
$e_{ijkN}$ $\mapsto 0$
$e_{ijkl}$ $\mapsto 0.$
Then it is a direct computation to sort these terms and put them together according to (2.13) for each basis element. The pairing of the resulting expressions (w.r.t. the tractor metric) with the $4$-tractor $T_{4}$ described explicitly in (2.8) then gives the mentioned quantities. ∎
The quantities for arbitrary curve satisfying $\Delta_{5}=0,\Delta_{4}\neq 0$ are given by the same formulae multiplied by $(-\Delta_{4})^{-\frac{1}{2}}$.
Example 2.1.
Quantities in dimension $3$. We have $5$ basis elements $e_{0}=e_{0123},e_{1}=e_{0124},e_{2}=e_{0134},e_{3}=e_{0234},e_{4}=e_{1234}$ and denote corresponding coordinates $w^{0},\dots,w^{4}$. The standard action (2.2) and tensoriality then give
$e_{0}\mapsto-X^{1}e_{3}+X^{2}e_{2}-X^{3}e_{1},\ \ \ e_{1}\mapsto X^{3}e_{4},\ \ \ e_{2}\mapsto-X^{2}e_{4},\ \ \ e_{3}\mapsto X^{1}e_{4},\ \ \ e_{4}\mapsto 0.$
In the coordinates $w^{a}$, we can write the corresponding transformation matrix $\rho_{4}(J)$ and thus $\exp(-\rho_{4}(J))$
$\rho_{4}(J)=\left(\begin{smallmatrix}0&0&0&0&0\\ -X^{3}&0&0&0&0\\ X^{2}&0&0&0&0\\ -X^{1}&0&0&0&0\\ 0&X^{3}&-X^{2}&X^{1}&0\end{smallmatrix}\right),\ \ \exp(-\rho_{4}(J))=\left(\begin{smallmatrix}1&0&0&0&0\\ -X^{3}&1&0&0&0\\ X^{2}&0&1&0&0\\ -X^{1}&0&0&1&0\\ -\frac{1}{2}\big{(}(X^{1})^{2}+(X^{2})^{2}+(X^{3})^{2}\big{)}&X^{3}&-X^{2}&X^{1}&1\end{smallmatrix}\right).$
Thus we describe elements of $\wedge^{4}\mathbb{R}^{5}$ in exponential coordinates $w^{a}$ as $\exp(-\rho_{4}(J))(w)$, and we get coordinates
$[\ w^{0},\ \ w^{0}X^{3}+w^{1},\ \ -w^{0}X^{2}+w^{2},\ \ w^{0}X^{1}+w^{3},$
$-\frac{1}{2}\big{(}(X^{1})^{2}+(X^{2})^{2}+(X^{3})^{2}\big{)}{w^{0}}-{w^{3}}{X^{1}}+{w^{2}}{X^{2}}-{w^{1}}{X^{3}}+{w^{4}}\ ]$
Now pairing these with $T_{4}$ with respect to the tractor metric gives conserved quantities. We get a generic family of quantities so that for each $a=0,\dots,4$, we choose $w^{a}=1$ and remaining coordinates zero.
We give the resulting quantities as follows. Below, the wedge product of three vectors is the scale given by the determinant of the corresponding matrix, and the wedge product of two vectors is the cross product. Then we get three quantities of type (1)
$Q^{0124}=u^{-4}X^{3}\cdot U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
$Q^{0134}=u^{-4}X^{2}\cdot U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
$Q^{0234}=u^{-4}X^{1}\cdot U\wedge A\wedge A^{\prime}-u^{-2}\>U\wedge A^{\prime}+3u^{-4}\>\langle U,A\rangle\>U\wedge A$
and we have one quantity of type (2)
$Q^{0123}=3u^{4}\langle U,A\rangle\>X\wedge U\wedge A-u^{-2}X\wedge U\wedge A^{\prime}+\frac{1}{2}|X|^{2}u^{-4}\>U\wedge A\wedge A^{\prime}.$
Let us note that in this dimension, the expression simplifies as the wedge product of four vectors vanishes. There is one quantity of type (3)
$Q^{1234}=-u^{-4}\>U\wedge A\wedge A^{\prime}$
and quantities of type (4) do not appear.
Example 2.2.
Logarithmic spirals. Let us consider a family of logarithmic spirals as follows
$\widehat{X}(t)=e^{t}\cos(ct)P_{0}+e^{t}\sin(ct)Q_{0}+R_{0},$ (2.14)
where $c$ is a constant, $P_{0},Q_{0},R_{0}$ are constant vectors, $\langle P_{0},Q_{0}\rangle=0$ and ${|P_{0}|=|Q_{0}|}.$ We compute quantities (2.9)-(2.12) for these curves. Direct computation gives
$U(t)$ $=e^{t}(\cos{(ct)}-\sin{(ct)})P_{0}+e^{t}(\cos{(ct)}+\sin{(ct)})Q_{0},$
$A(t)$ $=e^{t}((1-c^{2})\cos{(ct)}-2c\sin{ct})P_{0}+e^{t}((1-c^{2})\sin{(ct)}+2c\cos{(ct)}Q_{0},$
$A^{\prime}(t)$ $=e^{t}((c^{3}-3c)\sin{(ct)}+(-3c^{2}+1)\cos{(ct)})P_{0}$
$-e^{t}((c^{3}-3c)\cos{(ct)}+(3c^{2}-1)\sin{(ct)})Q_{0}.$
Then we compute
$u^{2}=e^{2t}(c^{2}+1)|P_{0}|^{2},\ \ \ \ \langle U,A\rangle=e^{2t}(c^{2}+1)|P_{0}|^{2},\ \ \ \ \langle U,A^{\prime}\rangle=e^{2t}(c^{4}-1)|P_{0}|^{2}$
thanks to the facts $|P_{0}|=|Q_{0}|$ and $\langle P_{0},Q_{0}\rangle=0$. Moreover, the expression $\epsilon^{ijk}(P_{0},Q_{0},R_{0})$ (up to the order of vectors) is the only non-vanishing sub-determinant of third order; other possibilities contain at least two same vectors and vanish. Analogously, expressions of the fourth order always vanish. Altogether, we compute
$Q^{0ijN}=\frac{c}{|P_{0}|^{2}}\epsilon^{ij}(P_{0},Q_{0}),\ \ \ \ Q^{0ijk}=\frac{c}{|P_{0}|^{2}}\epsilon^{ijk}(P_{0},Q_{0},R_{0}),\ \ \ \ Q^{ijkN}=Q^{ijkl}=0.$
In particular, the quantities are constant along spirals.
Moreover, we can find explicitly that along the curves $\widehat{X}(t)$, the tractor $\mathbb{A}$ takes form
$\mathbb{A}=\frac{e^{-t}}{|P_{0}|\sqrt{c^{2}+1}}e_{0}-\frac{\sqrt{c^{2}+1}(\cos{(ct)}P_{0}^{i}+\sin{(ct)}Q_{0}^{i})}{|P_{0}|}e_{i}-e^{t}|P_{0}|\sqrt{c^{2}+1}e_{N}.$
Then the square of the length of this tractor for the tractor metric is $\alpha_{1}=c^{2}-1$. Similarly, one computes the tractor $\mathbb{A^{\prime}}$ and the square of its length $\alpha_{2}=c^{4}-c^{2}+1$. In particular,
$\Delta_{4}=\alpha_{1}^{2}-\alpha_{2}=-c^{2}.$
It is a more nasty but analogous computation to find that $\Delta_{5}=0$.
Let us note that one of the possible definitions of logarithmic spirals is via invariants, [14]. We say that the curve is a spiral if $\kappa_{1}$ is constant and $\Delta_{5}=0$, where
$\kappa_{1}=-\frac{1}{2}(-\Delta_{4})^{-\frac{5}{2}}(\alpha_{1}\Delta_{4}^{2}-\frac{1}{2}\Delta_{4}\Delta^{\prime\prime}_{4}+\frac{9}{16}(\Delta^{\prime}_{4})^{2}).$
In fact, if we assume $\Delta_{4}$ constant, we also get $\alpha_{1}$ constant from the fact that $\kappa_{1}$ constant. The condition $\alpha_{1}$ constant gives a special parametrization which generalizes projective parametrization; the projective parametrization corresponds to $\alpha_{1}=0$. In general, if $\Delta_{4}$ and $\alpha_{1}$ are constant, then $\kappa_{1}$ is constant, but the opposite direction does not necessarily hold in general. A different viewpoint on spirals was introduced in [7], where the author focuses on homogeneous curves and studies their symmetry algebras.
3. Conserved quantities for the conformal Mercator equation
Here we come to our second generalization of conformal circles, that is, the class of parametrized curves which are solutions of the conformal Mercator equation, and present their conserved quantities. Note that the logarithmic spiral (2.14) is also a solution [5]. Thus the generalization here and the one in Section 2.2 describe some common curves.
3.1. Conformal Mercator equation in the conformally-flat case
Choosing the Euclidean metric $\delta$ from the conformal class, the conformal Mercator equation [5] is given by
$\frac{dC}{dt}=0,\;C=u^{-2}\left(A^{\prime}-u^{-2}|A|^{2}U-2u^{-2}\langle A,U\rangle A+4u^{-4}\langle A,U\rangle^{2}U-2u^{-2}\langle A^{\prime},U\rangle U\right),$ (3.15)
where $U$ and $A$ are respectively the velocity and acceleration vector fields of a curve $\gamma(t),$ and $A^{\prime}=\frac{dA}{dt},$ as before.
Equation (3.15) is the Euler-Lagrange equation of the Lagrangian [5]
$L=\frac{d}{dt}\left(u^{-2}\langle U,A\rangle\right)+\frac{1}{2}u^{-2}|A|^{2}-u^{-4}\langle U,A\rangle^{2},$ (3.16)
where the variation and its derivatives up to second order are assumed to vanish at end points. This fact enables us to apply Noether’s principle to find conserved quantities (or first integrals - functions that are constant on all solutions) of (3.15).
3.2. Conserved quantities from conformal Killing vector fields
Our derivation of conserved quantities is based on a simplest version of Noether’s theorem, which we shall describe below. (See e.g. [1].) Suppose an ODE is the Euler-Lagrange equation of a functional
[…]
meaning that all solutions of the ODE are critical curves of $I(\gamma).$ A one-parameter group of symmetries of the Lagrangian $L$ is a one-parameter group of transformations $\{\phi_{s}\}$ that preserves $L$ in the sense that
$L(\gamma(t))=L(\phi_{s}(\gamma(t))),$ (3.17)
where $L(\gamma(t))$ denotes the Lagrangian evaluated on a curve $\gamma,$ while $L(\phi_{s}(\gamma(t)))$ denotes the one evaluated on the transformed curve $\phi_{s}(\gamma(t)).$ Noether’s theorem states that there exists an associated conserved quantity for solutions of the Euler-Lagrange equation for each such symmetry group. [^3] The general statement of Noether’s theorem can be found in e.g. [13], where the functional is a function on a set of $k$-dimensional submanifolds of a manifold $M,$ $1\leq k<{\rm dim}\,M,$ generalizing a set of curves in $M,$ and leading to the Euler-Lagrange equation as a PDE. In addition, there exists a conserved quantity for each symmetry group of the functional, which is more general than a symmetry group of the Lagrangian. Note that if $\{\phi_{s}\}$ is a symmetry group of the Lagrangian, then it is also a symmetry group of the Euler-Lagrange equation, which means that if $\gamma(t)$ is a solution of the ODE, then $\phi_{s}(\gamma(t))$ is also a solution.
Based on Noether’s principle, let us now derive an expression for the conserved quantities for solutions of the conformal Mercator equation (3.15). Let $\gamma$ be a solution of (3.15) and $\{\phi_{s}\}$ a one-parameter group of symmetries of the Lagrangian (3.16). Consider a one-parameter family of solution curves $\Gamma:[s_{0},s_{1}]\times[t_{0},t_{1}]\to\mathbb{R}^{n}$ given by $\Gamma(s,t)=\phi_{s}(\gamma(t)).$ Choose the parameter $s$ such that $0\in[s_{0},s_{1}]$ and $\phi_{0}={\rm Id},$ i.e. $\Gamma(0,t)=\gamma(t).$ If we restrict the range of the parameter $t$ so that the generator of $\{\phi_{s}\}$ is nowhere tangent to $\gamma,$ then for a sufficiently small range of $s$ around $0,$ the family $\Gamma$ is a $2$-dimensional submanifold of $\mathbb{R}^{n}.$
Now, let $U$ (where we abuse the notation) denote the vector field on $\Gamma$ whose restriction to each solution curve $\phi_{s}(\gamma(t))$ ($s$ fixed) is the velocity of the curve. Also, let $V$ be the generator of $\{\phi_{s}\}$ restricted to $\Gamma.$ By definition, $U$ is invariant under the flow of $V,$ and thus $[U,V]=0.$
Next, consider the Lagrangian function on $\Gamma$ given by
$L(s,t):=L(\phi_{s}(\gamma(t)))=\nabla_{U}\left(u^{-2}\langle U,A\rangle\right)+\frac{1}{2}u^{-2}|A|^{2}-u^{-4}\langle U,A\rangle^{2},$
where $\nabla$ is the Levi-Civita connection of $\delta$ and $\nabla_{U}$ the covariant derivative along each curve $\phi_{s}(\gamma(t)).$ Equation (3.17) implies that $V(L(s,t))=0.$ Then, by direct calculation with some rearrangements, it can be shown that
$0=V(L)=\nabla_{U}F+\langle\nabla_{U}C,V\rangle,$ (3.18)
where $F$ is a function on $\Gamma$ given by
$F=\nabla_{U}\langle W,\nabla_{U}V\rangle+\langle\nabla_{U}W,\nabla_{U}V\rangle-\langle C,V\rangle,\quad\mbox{with}\quad W=u^{-2}U,$
and $C$ is defined in (3.15). Let us note that in obtaining (3.18), we have used the flatness of $\delta$ and the fact that $[U,V]=0.$ These imply that $\nabla_{V}U=\nabla_{U}V$ and $\nabla_{V}\nabla_{U}U=\nabla_{U}\nabla_{V}U.$
Now restricting (3.18) to each solution curve $\phi_{s}(\gamma(t))$ which satisfies (3.15), $\nabla_{U}C=0,$ we have that $\nabla_{U}F=0$ upon restriction. Hence $F$ is constant on each solution.
The proposition below follows from the fact that the Lagrangian (3.16) is invariant under a one-parameter group of conformal transformations. [^4] One can verify directly that the Lagrangian (3.16) satisfies (3.17) for any one-parameter group conformal transformations $\phi_{s}.$
Proposition 3.1.
Let $\gamma:t\mapsto\gamma(t)$ be a smooth curve, and $U$ and $A$ be its velocity and acceleration vector fields, respectively. Let $V$ be a conformal Killing vector field restricted to $\gamma.$ Then the function
$F=\frac{d}{dt}\langle W,V^{\prime}\rangle+\langle W^{\prime},V^{\prime}\rangle-\langle C,V\rangle,\qquad W=u^{-2}U,$ (3.19)
where $C$ is given in (3.15) and $\,{}^{\prime}$ denotes $\frac{d}{dt},$ is constant on any curve $\gamma$ that satisfies the conformal Mercator equation (3.15).
Proof.
The proposition follows from the derivation based on Noether’s principle as discussed above. However, it can also be proved by directly calculating $\nabla_{U}F=F^{\prime}$ and show that it vanishes if the conformal Mercator equation and the conformal Killing equation hold. Let us outline the calculation here briefly.
First recall that a conformal Killing vector field $V$ satisfies the conformal Killing equation
$\nabla_{(i}V_{j)}=\lambda\,\delta_{ij},\qquad\lambda=\frac{1}{n}\nabla_{i}V^{i},$ (3.20)
where we use the Einstein summation convention. This implies that upon restriction to a curve $\gamma,$ the vector field $V$ satisfies
$\langle\nabla_{U}V,U\rangle=\lambda\,u^{2},\quad\mbox{i.e.}\quad\langle\nabla_{U}V,W\rangle=\lambda.$ (3.21)
Now rewriting $F$ as
$F=\nabla_{U}\langle W,\nabla_{U}V\rangle-2\langle W,A\rangle\langle W,\nabla_{U}V\rangle+\langle u^{-2}A,\nabla_{U}V\rangle-\langle C,V\rangle$
gives
$\nabla_{U}F$ $=$ $\nabla_{U}^{2}\langle W,\nabla_{U}V\rangle-2\langle W,A\rangle\nabla_{U}\langle W,\nabla_{U}V\rangle-u^{-2}|A|^{2}\langle W,\nabla_{U}V\rangle$
$+\langle u^{-2}A,\nabla_{U}^{2}V\rangle-\langle\nabla_{U}C,V\rangle,$
where the last term vanishes by the conformal Mercator equation $\nabla_{U}C=0.$
Using the conformal Killing equation (3.21), we have that
$\nabla_{U}F=\nabla_{U}^{2}\lambda-2\langle W,A\rangle\nabla_{U}\lambda-u^{-2}|A|^{2}\lambda+\langle u^{-2}A,\nabla_{U}^{2}V\rangle.$ (3.22)
Then the last term of (3.22) can be computed using a differential consequence of the conformal Killing equation (3.20) and the fact that $\nabla$ is flat. This results in
$\nabla_{U}F=\nabla_{U}^{2}\lambda-d\lambda(A)=U^{i}U^{j}\nabla_{i}\nabla_{j}\lambda.$
Now, if $V$ is Killing, then $\lambda$ is zero. Also, we have that $\lambda=\frac{1}{n}\nabla_{i}V^{i}$ is constant for the generator of dilatation, and is linear in spacetime coordinates for those of special conformal transformations. It follows that $\nabla_{U}F=0$ for all conformal Killing vector fields. ∎
3.3. Expressions for conserved quantities and examples
The conformal Killing vector fields can be divided into four types, namely the generators of translation, rotation, dilatation and special conformal transformation. Using the same notation as in Proposition 3.1, below we give an expression for the conserved quantity (3.19) for a conformal Killing vector field $V$ of each type. Note that tautologically $C$ is a constant vector field along any solution curve of (3.15). Also, the first term of (3.19) vanishes for all conformal Killing vector fields $V$ except for the generator of a special conformal transformation.
Proposition 3.2.
Let $(x^{i})$ be the standard coordinates for $\mathbb{R}^{n},$ and $\gamma$ be a solution to the conformal Mercator equation (3.15) given by $(x^{i})=(\gamma^{i}(t))=:(X^{i}(t)).$
Let $V_{T},$ $V_{R},$ $V_{D}$ and $V_{S}$ be the generators of translation, rotation, dilatation and special conformal transformation, respectively, given by
$V_{T}=T^{i}\frac{\partial}{\partial x^{i}},\quad V_{R}={R_{j}}^{i}x^{j}\frac{\partial}{\partial x^{i}},\quad V_{D}=a\,x^{i}\frac{\partial}{\partial x^{i}},\quad V_{S}=\left(x_{j}x^{j}S^{i}-2S_{j}x^{j}x^{i}\right)\frac{\partial}{\partial x^{i}},$
where $T^{i},{R_{i}}^{j},a,S^{i}$ are constant and arbitrary, with $R_{ij}=-R_{ji},$ and the indices are raised and lowered by $\delta.$ Also, let $T:=T^{i}\frac{\partial}{\partial x^{i}}$ and $S:=S^{i}\frac{\partial}{\partial x^{i}}.$
Then the corresponding conserved quantities, $F$ (3.19), are given by
$F_{T}$ $=$ $-\langle C,T\rangle$ (3.23)
$F_{R}$ $=$ $R_{ij}\left(u^{-2}U^{i}A^{j}+C^{i}X^{j}\right)$ (3.24)
$F_{D}$ $=$ $-a\left(u^{-2}\langle U,A\rangle+\langle C,X\rangle\right)$ (3.25)
$F_{S}$ $=$ $2\langle S,{\mathcal{Y}}\rangle\quad\mbox{with}$ (3.26)
${\mathcal{Y}}$ $=$ $u^{-2}\langle U,X\rangle A-\left(1+u^{-2}\langle A,X\rangle\right)U\left(u^{-2}\langle U,A\rangle+\langle C,X\rangle\right)X-\frac{|X|^{2}}{2}C,$
where we use $X$ to denote the vector field $X^{i}\frac{\partial}{\partial x^{i}}$ along $\gamma.$
Proof.
The proposition follows by direct substitution of $V_{T},$ $V_{R},$ $V_{D},$ $V_{S}$ into (3.19), where upon restriction to $\gamma$ we have $V_{R}={R_{j}}^{i}X^{j}\frac{\partial}{\partial x^{i}},$ $V_{D}=a\,X$ and $V_{S}=|X|^{2}S-2\langle S,X\rangle X.$
Note that for translations, choosing $T$ to be each coordinate vector field $T=\frac{\partial}{\partial x^{i}},$ the quantity $F_{T}=-C^{i}$ is just minus each component of the tautologically constant vector field $C.$
Example 3.1.
It was shown in [5] that the logarithmic spiral (2.14),
$\widehat{X}(t)=e^{t}\cos(ct)P_{0}+e^{t}\sin(ct)Q_{0}+R_{0},$
is the general solution of the reduction of (3.15) under the condition $C=0.$ For this solution, the conserved quantities (3.23) - (3.26) are given by
$F_{T}=0,\quad F_{R}=\frac{c}{|P_{0}|^{2}}R_{kl}P_{0}^{k}Q_{0}^{l},\quad F_{D}=-a,\quad F_{S}=2\langle S,-c\frac{\langle Q_{0},R_{0}\rangle}{|P_{0}|^{2}}P_{0}+c\frac{\langle P_{0},R_{0}\rangle}{|P_{0}|^{2}}Q_{0}+R_{0}\rangle.$
Let us note the similarity between the quantity $Q^{0ijN}=\frac{c}{|P_{0}|^{2}}\epsilon^{ij}(P_{0},Q_{0})$ of Example 2.2 and $F_{R}$ here. In fact, for a fixed $i<j,$ choosing the anti-symmetric matrix $R$ to be such that $R_{ij}=1,$ $R_{ji}=-1$ and all other entries vanish, one has $F_{R}=Q^{0ijN}.$ We will come back to this example again in Section 4 after we establish the relation between the conserved quantities of Proposition 2.3 and Proposition 3.2.
Example 3.2.
Recall that under the stereographic projection, the inverse image of logarithmic spirals on a plane are loxodromes on $S^{2},$ which are curves that cut all meridians at a constant angle. Let us consider the logarithmic spiral (2.14) in two dimensions.
Without loss of generality, one can choose $P_{0}=(1,0),$ $Q_{0}=(0,1)$ and $R_{0}=(0,0).$ Then the inverse image of the curve is a loxodrome which cuts all meridians at the angle $\cos^{-1}(1/\sqrt{1+c^{2}}).$ The conserved quantity which corresponds to this constant angle is $F_{R}$; choosing $R=\left(\begin{smallmatrix}0&1\\ -1&0\end{smallmatrix}\right),$ we have that $F_{R}=c.$
Example 3.3.
Due to conformal invariance of the conformal Mercator equation (3.15), the special conformal transformation of (2.14),
$X(t)=\frac{\widehat{X}(t)-|\widehat{X}(t)|^{2}B}{1-2\langle\widehat{X}(t),B\rangle+|B|^{2}|\widehat{X}(t)|^{2}},$ (3.27)
where $B$ is a constant vector, is also a solution of (3.15) [5].
Below we give the values of the conserved quantities (3.23) - (3.26), evaluated on (3.27):
$F_{T}$ $=$ $-\langle C,T\rangle,\qquad\mbox{where}$
$C$ $=$ $\frac{2}{|P_{0}|^{2}}\Big{(}c\langle Q_{0},|B|^{2}R_{0}-B\rangle P_{0}-c\langle P_{0},|B|^{2}R_{0}-B\rangle Q_{0}-|B|^{2}|P_{0}|^{2}R_{0}$
$\qquad+\big{(}2c\langle P_{0},R_{0}\rangle\langle Q_{0},B\rangle-2c\langle Q_{0},R_{0}\rangle\langle P_{0},B\rangle+(2\langle R_{0},B\rangle-1)|P_{0}|^{2}\big{)}\,B\Big{)}$
$F_{R}$ $=$ $\frac{c}{|P_{0}|^{2}}[P_{0},Q_{0}]_{R}-2[B,R_{0}]_{R}+\frac{2c}{|P_{0}|^{2}}\Big{(}[Q_{0},B]_{R}\langle P_{0},R_{0}\rangle-[P_{0},B]_{R}\langle Q_{0},R_{0}\rangle\Big{)},$
$\mbox{where}\quad[P_{0},Q_{0}]_{R}:=R_{ij}P_{0}^{i}Q_{0}^{j},\;\mbox{etc.}$.
$F_{D}$ $=$ $-a\left(1-2\langle R_{0},B\rangle+\frac{2c}{|P_{0}|^{2}}\big{(}\langle P_{0},B\rangle\langle Q_{0},R_{0}\rangle-\langle Q_{0},B\rangle\langle P_{0},R_{0}\rangle\big{)}\right)$
$F_{S}$ $=$ $2\langle S,{\mathcal{Y}}\rangle,\quad\mbox{where}\quad{\mathcal{Y}}=-c\frac{\langle Q_{0},R_{0}\rangle}{|P_{0}|^{2}}P_{0}+c\frac{\langle P_{0},R_{0}\rangle}{|P_{0}|^{2}}Q_{0}+R_{0}.$
Let us end this section by noting that, as conformal circles are also solutions of the conformal Mercator equation (3.15), the function $F$ (3.19) also gives a conserved quantity for conformal circles for each conformal Killing vector field $V.$ We summarize this result in Appendix B.
4. Hamiltonian formulation and relation between conserved quantities
In this section we write our conserved quantities in terms of the Hamiltonian phase space variables associated to the conformal Mercator equation. The advantage of this formulation is that the expressions for the quantities are all simplified to polynomials. This helps us to notice and write down relations between them. In particular, we manage to write down the conserved quantities of Proposition 2.3 in terms of the basis quantities for (3.23) - (3.26).
4.1. Hamiltonian formalism
Here we briefly review the Hamiltonian formulation of the conformal Mercator equation (3.15) given in [5]. Using the Ostrogradsky construction to the second order derivative Lagrangian [2]
$L_{1}=\frac{1}{2}u^{-2}|A|^{2}-u^{-4}\langle U,A\rangle^{2},$
which is the Lagrangian $L$ (3.16) without the first term [^5] Recall that the conformal Mercator equation (3.15) is the Euler-Lagrange equation of both Lagrangains $L$ and $L_{1},$ as the first term of $L$ (3.16) is integrated to a boundary term. In Section 3, we use $L$ as we require conformal invariance. Here we simply describe a Hamiltonian system which is equivalent to (3.15)., the corresponding Hamiltonian phase space is of $4n$ dimensions, with $X^{i}$ and $U^{i},$ $i=1,\dots,n,$ as the first $2n$ canonical coordinates. The other $2n$ coordinates are given by the momenta $\mathcal{P}=(\mathcal{P}_{i})$ and $\mathcal{R}=(\mathcal{R}_{i})$ conjugate to $X$ and $U,$ respectively, which are defined by
$\mathcal{P}_{i}=\frac{\partial L_{1}}{\partial U^{i}}-\frac{d}{dt}\frac{\partial L_{1}}{\partial A^{i}}\qquad\mbox{and}\qquad\mathcal{R}_{i}=\frac{\partial L_{1}}{\partial A^{i}}.$
It turns out that
$\begin{split}\mathcal{P}&=-u^{-2}\left(A^{\prime}-u^{-2}|A|^{2}U-2u^{-2}\langle A,U\rangle A+4u^{-4}\langle A,U\rangle^{2}U-2u^{-2}\langle A^{\prime},U\rangle U\right)\\ &=-C,\end{split}$ (4.28)
$\mathcal{R}$ $=u^{-2}A-2u^{-4}\langle U,A\rangle U.$ (4.29)
The Hamiltonian is then given by the Legendre transform ${H=\langle\mathcal{P},X\rangle+\langle\mathcal{R},U\rangle-L_{1},}$ which gives
$H\;=\;\langle\mathcal{P},U\rangle-\langle\mathcal{R},U\rangle^{2}+\frac{1}{2}u^{2}|\mathcal{R}|^{2}.$ (4.30)
It follows that the conformal Mercator equation (3.15) is equivalent to the $4n$ first order Hamilton’s equations
${X^{i}}^{\prime}=\frac{\partial H}{\partial\mathcal{P}_{i}},\qquad{U^{i}}^{\prime}=\frac{\partial H}{\partial\mathcal{R}_{i}},\qquad{\mathcal{P}_{i}}^{\prime}=-\frac{\partial H}{\partial X^{i}},\qquad{\mathcal{R}_{i}}^{\prime}=-\frac{\partial H}{\partial U^{i}},$
that is, [5]
$X^{\prime}=U,\quad U^{\prime}=u^{2}{\mathcal{R}}-2\langle U,\mathcal{R}\rangle U,\quad{\mathcal{P}}^{\prime}=0,\quad{\mathcal{R}}^{\prime}=-|\mathcal{R}|^{2}U+2\langle U,\mathcal{R}\rangle\mathcal{R}-\mathcal{P}.$ (4.31)
Let us recall (see e.g. [1]) that given a function $f(X,U,{\mathcal{P}},{\mathcal{R}})$ on the Hamiltonian phase space, the derivative of $f$ along a solution curve of (4.31) is given by
$\frac{df}{dt}=\{f,H\},$
[…]
$\{f,g\}=\frac{\partial f}{\partial X^{i}}\frac{\partial g}{\partial\mathcal{P}_{i}}+\frac{\partial f}{\partial U^{i}}\frac{\partial g}{\partial\mathcal{R}_{i}}-\frac{\partial g}{\partial X^{i}}\frac{\partial f}{\partial\mathcal{P}_{i}}-\frac{\partial g}{\partial U^{i}}\frac{\partial f}{\partial\mathcal{R}_{i}},$
using the summation convention.
It then follows that a function $f(X,U,{\mathcal{P}},{\mathcal{R}})$ is constant along a solution curve if and only if $\{f,H\}=0.$
4.2. Conserved quantities in phase-space variables
It turns out that the conserved quantities for solutions of the conformal Mercator equation, (3.23) - (3.26), are all polynomials in the phase space variables. Treating $X,U,{\mathcal{P}},{\mathcal{R}}$ as $n$-vectors, the quantities are given by
$F_{T}$ $=$ $\langle T,\mathcal{P}\rangle,$
$F_{R}$ $=$ $R_{ij}(X^{i}\mathcal{P}^{j}+U^{i}\mathcal{R}^{j}),$
$F_{D}$ $=$ $a\,(\langle X,\mathcal{P}\rangle+\langle U,\mathcal{R}\rangle),$
$F_{S}$ $=$ $\langle S,|X|^{2}\mathcal{P}+2\langle X,U\rangle\mathcal{R}-2(\langle X,\mathcal{P}\rangle+\langle U,\mathcal{R}\rangle)X-2(1+\langle X,\mathcal{R}\rangle)U\rangle,$
where in obtaining these expressions we have used the conformal Mercator equation in Hamilton’s form (4.31), and also the relations (4.28), (4.29), which imply
$A$ $=$ $-2\langle\mathcal{R},U\rangle U+u^{2}\mathcal{R},$
$A^{\prime}$ $=$ $\left(2\langle U,\mathcal{P}\rangle+4\langle U,\mathcal{R}\rangle^{2}-u^{2}|\mathcal{R}|^{2}\right)U-2u^{2}\langle U,\mathcal{R}\rangle\mathcal{R}-u^{2}\mathcal{P}.$
Among the above quantities, we find that there are precisely $(n+1)(n+2)/2$ functionally independent conserved quantities, corresponding to the basis vector fields of the Lie algebra of the conformal Killing vector fields. Moreover, the Poisson bracket relations of these basis quantities follow the Lie bracket relations of the associated basis vector fields in an analogous way.
In writing down the relation with the conserved quantities of Proposition 2.3, we shall use the following $(n+1)(n+2)/2$ basis quantities, which correspond to a choice of basis conformal Killing vector fields:
$E_{T}^{i}$ $=$ $\mathcal{P}^{i}$ (4.32)
$E_{R}^{ij}$ $=$ $\epsilon^{ij}(X,\mathcal{P})+\epsilon^{ij}(U,\mathcal{R})\qquad\quad$ (4.33)
$E_{D}$ $=$ $\langle X,\mathcal{P}\rangle+\langle U,\mathcal{R}\rangle$ (4.34)
$E_{S}^{i}$ $=$ $|X|^{2}\mathcal{P}^{i}+2\langle X,U\rangle\mathcal{R}^{i}-2(\langle X,\mathcal{P}\rangle+\langle U,\mathcal{R}\rangle)X^{i}-2(1+\langle X,\mathcal{R}\rangle)U^{i},\qquad\quad$ (4.35)
where $i,j=1,\dots,n,$ and $\epsilon^{ij}$ is the elementary alternating tensor of rank $2,$ defined in Section 2.3.
Similarly, the conserved quantities (2.9) - (2.12) of Proposition 2.3 can be written as
$Q^{0ijN}$ $=-\langle\mathcal{R},U\rangle\epsilon^{ij}(U,\mathcal{R})+\epsilon^{ij}(U,\mathcal{P})-\epsilon^{ijk}(U,\mathcal{R},\mathcal{P})X_{k}$ (4.36)
$\begin{split}Q^{0ijk}&=-\Big{(}\,\langle U,\mathcal{R}\rangle\epsilon^{ijk}(X,U,\mathcal{R})+\epsilon^{ijk}(U,X,\mathcal{P})\\ &\qquad\quad+\frac{1}{2}|X|^{2}\epsilon^{ijk}(U,\mathcal{R},\mathcal{P})+\epsilon^{ijkl}(X,U,\mathcal{R},\mathcal{P})X_{l}\,\Big{)}\quad\qquad\end{split}$ (4.37)
$Q^{ijkN}$ $=-\epsilon^{ijk}(U,\mathcal{R},\mathcal{P})$ (4.38)
$Q^{ijkl}$ $=-\epsilon^{ijkl}(X,U,\mathcal{R},\mathcal{P}).$ (4.39)
4.3. Relation between conserved quantities
In the following theorem, we write the conserved quantities of Proposition 2.3, (4.36)-(4.39), in terms of the quantities (4.32)-(4.35). For convenience, let $E_{T}:=(\,E_{T}^{i}\,)$ and $E_{S}:=(\,E_{S}^{i}\,)$ denote the $n$-vectors whose components are $E_{T}^{i}$ and $E_{S}^{i},$ respectively.
Theorem 4.1.
With the notations defined above, the following identities hold.
$Q^{0ijN}$ $=$ $\frac{1}{2}\epsilon^{ij}(E_{T},E_{S})-E_{D}\,E_{R}^{ij}$ (4.40)
$Q^{0ijk}$ $=$ $\frac{3}{2}E_{R}^{[ij}\,E_{S}^{k]}$ (4.41)
$Q^{ijkN}$ $=$ $-3\,E_{R}^{[ij}\,E_{T}^{k]}$ (4.42)
$Q^{ijkl}$ $=$ $\frac{6}{E_{D}}\,E_{R}^{[ij}\,E_{S}^{k}\,E_{T}^{l]}.$ (4.43)
Proof.
The proof is by direct calculation, where we use identities from the property of the determinant such as $\epsilon^{ijk}(U,\mathcal{R},\mathcal{P})=3\,\epsilon^{[ij}(U,\mathcal{R})\,\mathcal{P}^{k]}.$ Note also that $Q^{0ijk}$ (4.37) can be written as
$Q^{0ijk}$ $=$ $\frac{1}{2}|X|^{2}\epsilon^{ijk}(U,\mathcal{R},\mathcal{P})+\langle X,U\rangle\epsilon^{ijk}(X,\mathcal{P},\mathcal{R})-(1+\langle X,\mathcal{R}\rangle)\epsilon^{ijk}(X,\mathcal{P},U)$
$-(\langle X,\mathcal{P}\rangle+\langle U,\mathcal{R}\rangle)\epsilon^{ijk}(U,\mathcal{R},X),$
using the equality
$\epsilon^{ijkl}(X,U,\mathcal{R},\mathcal{P})X_{l}$ $=$ $-|X|^{2}\epsilon^{ijk}(U,\mathcal{R},\mathcal{P})+\langle X,U\rangle\epsilon^{ijk}(X,\mathcal{R},\mathcal{P})-\langle X,\mathcal{R}\rangle\epsilon^{ijk}(X,U,\mathcal{P})$
$+\langle X,\mathcal{P}\rangle\epsilon^{ijk}(X,U,\mathcal{R}),$
which also follows from the property of the determinant. ∎
Note that although being proved here in the Hamiltonian formulation, the identities of Theorem 4.1 hold regardless of whether the quantities are expressed in the phase space or ‘spacetime’ variables.
Example 4.1.
Here we come back to the logarithmic spiral (2.14),
$\widehat{X}(t)=e^{t}\cos(ct)P_{0}+e^{t}\sin(ct)Q_{0}+R_{0}.$
It can be shown that
$E_{T}^{i}=0,E_{R}^{ij}=\frac{c}{|P_{0}|^{2}}\epsilon^{ij}(P_{0},Q_{0}),\quad E_{D}=-1,\quad E_{S}^{i}=-2c\frac{\langle Q_{0},R_{0}\rangle}{|P_{0}|^{2}}P_{0}^{i}+2c\frac{\langle P_{0},R_{0}\rangle}{|P_{0}|^{2}}Q_{0}^{i}+2R_{0}^{i}.$
Then one can use Theorem 4.1 to compute the conserved quantities of Proposition 2.3. Since $E_{T}^{i}=0,$ we have that $Q^{ijkN}=Q^{ijkl}=0.$ Then
$Q^{0ijN}=-E_{D}\,E_{R}^{ij}=\frac{c}{|P_{0}|^{2}}\epsilon^{ij}(P_{0},Q_{0}),$
and
$Q^{0ijk}=\frac{3}{2}E_{R}^{[ij}\,E_{S}^{k]}=3\frac{c}{|P_{0}|^{2}}\epsilon^{[ij}(P_{0},Q_{0})\,R_{0}^{k]}=\frac{c}{|P_{0}|^{2}}\epsilon^{ijk}(P_{0},Q_{0},R_{0}).$
Cf. Example 2.2.
Example 4.2.
In $3$ dimensions, there are $3$ quantities for $Q^{0ij4};$ let $Q_{1}$ denote a $3$-vector given by $Q_{1}=(Q^{0234},-Q^{0134},Q^{0124}).$ There is only one quantity for each of $Q^{0ijk}$ and $Q^{ijk4}$, let $Q_{2}:=Q^{0123}$ and $Q_{3}:=Q^{1234}.$ Also recall that $Q^{ijkl}$ does not appear in $3$ dimensions.
[…]
$Q_{1}$ $=$ $\frac{1}{2}E_{T}\wedge E_{S}-E_{D}E_{R},$
$Q_{2}$ $=$ $\frac{1}{2}\langle E_{R},E_{S}\rangle,$
$Q_{3}$ $=$ $-\langle E_{T},E_{R}\rangle.$
In addition, the Hamiltonian (4.30), which is also a conserved quantity can be written as
$H=\frac{1}{2}\left(|E_{R}|^{2}-\langle E_{T},E_{S}\rangle-{(E_{D})}^{2}\right).$
As an aside, let us note that to confirm Liouville integrability of the conformal Mercator equation (4.31) in $3$ dimensions, one needs to show that there exists a set of $6$ involutive independent conserved quantities. Recall that two functions $f,g$ on a Hamiltonian phase space is said to be involutive if $\{f,g\}=0.$ However, so far we are unable to generate such a set from the $16$ conserved quantities ($E_{T},$ $E_{R},$ $E_{D},$ $E_{S},$ $Q_{1},$ $Q_{2},$ $Q_{3},$ $H$). The largest involutive independent set we have found consists of $5$ conserved quantities, that is $\{E_{T}^{i},Q_{3},H\},$ $i=1,2,3.$
Acknowledgements
The authors would like to thank Maciej Dunajski for introducing them to the problem. They are also grateful to Maciej Dunajski, Jan Gregorovič, Rod Gover, Wojciech Kryński, Josef Šilhan and Vojtěch Žádník for helpful discussions. The authors would also like to thank the Isaac Newton Institute for Mathematical Sciences, Cambridge, for support and hospitality during the programme Twistor theory, where some research on this paper was carried over. This work was supported by EPSRC grant EP/Z000580/1. L.Z. is supported by the grant GACR 24-10887S Cartan supergeometries and Higher Cartan geometries.
Appendix A Conserved quantities for conformal circles from parallel $3$-tractor
Let us give here for completeness the quantities for conformal circles coming from pairing of the $3$-tractor $\mathbb{T}\wedge\mathbb{U}\wedge\mathbb{A}$ and parallel tractors corresponding to CKY 2-forms aka elements of $\wedge^{3}\mathbb{R}^{n+2}$. Computations analogous to Section 2.3 give that the $3$-tractor takes form
$T_{3}=u^{-3}U^{[i}A^{j]}\>e_{0ij}-u^{-1}U^{i}\>e_{0iN}$
where we use Einstein summation convention, $[ab..c]$ denotes the anti-symmetrization over indices in the bracket and $i,j=1,\dots,n$. One can see that this corresponds to the results presented in [Section 4.4.3.][9]. Using notation analogous to Section 2.3, we get the quantities as follows:
$Q^{0iN}$ $=u^{-1}U^{i}+u^{-3}\epsilon^{ik}(U,A)X_{k}$
$Q^{0ij}$ $=-u^{-1}\epsilon^{ij}(X,U)+\frac{1}{2}u^{-3}|X|^{2}\epsilon^{ij}(U,A)-u^{-3}\epsilon^{ijk}(X,U,A)X_{k}$
$Q^{0ij}$ $=u^{-3}\epsilon^{ij}(U,A)$
$Q^{ijk}$ $=u^{-3}\epsilon^{ijk}(X,U,A).$
Appendix B Conserved quantities for conformal circles from Lagrangian formulation
It was shown in [5] that conformal circles correspond to the class of solutions of the conformal Mercator equation (3.15) satisfying the condition
$C=\left(\frac{1}{2}u^{-4}|A|^{2}-2u^{-6}\langle A,U\rangle^{2}\right)U+u^{-4}\langle A,U\rangle A.$
Then (3.15) reduces to the conformal circle equation
$A^{\prime}-3u^{-2}\langle A,U\rangle A-\frac{3}{2}u^{-2}|A|^{2}U=0,$ (B.44)
the general solution of which is a projectively parametrized circle
$X(t)=X_{0}+\frac{tU_{0}+t^{2}A_{0}}{1+t^{2}|A_{0}|^{2}},$
where $X_{0},U_{0},A_{0}$ are constant vectors, $|U_{0}|=1$ and $\langle U_{0},A_{0}\rangle=0.$
It turns out that for conformal circles, the conserved quantity of Proposition 3.1 takes a simple form. To see this, first note that the function $F$ (3.19) can be written as
$F=\frac{d^{2}}{dt^{2}}\langle W,V\rangle-\langle Z,V\rangle,\quad\mbox{where}\quad Z=\frac{1}{2}u^{-2}A^{\prime}+\frac{3}{2}W^{\prime\prime}-\langle W,A^{\prime}\rangle W.$
Now, by virtue of (B.44), it can be shown that the vector field $Z$ vanishes on any conformal circle. This gives
$F=\frac{d^{2}}{dt^{2}}\langle W,V\rangle.$
Recall that for all conformal Killing vector fields $V$ except for the generator of a special conformal transformation, we have that $\langle W,V^{\prime}\rangle$ is constant, and thus $F=\frac{d}{dt}\langle W^{\prime},V\rangle.$
Appendix C Note on tractor equation
The general principle states, that the middle aka invariant slot of the tractor equation
$\mathbb{A}^{\prime\prime}=\frac{\Delta^{\prime}_{4}}{2\Delta_{4}}\mathbb{A}^{\prime}-\alpha_{1}\mathbb{A}+\frac{1}{2\Delta_{4}}(\alpha^{\prime}_{1}(2\alpha_{2}-\Delta_{4})-\alpha_{1}\alpha^{\prime}_{2})\mathbb{U}+$
$\frac{1}{4\Delta_{4}}(2\alpha_{1}(\alpha^{\prime}_{1})^{2}-4\Delta_{4}^{2}-2\alpha^{\prime\prime}_{2}\Delta_{4}-\alpha^{\prime}_{1}\alpha^{\prime}_{2})\mathbb{T}$
gives the equation for the corresponding curves satisfying $\Delta_{5}=0$. This can be done in a general setting but will be nasty. Involving the conditions $\Delta^{\prime}_{4}=0$ and $\alpha^{\prime}_{1}=0$, the middle slot gives an equation of the form
$\begin{split}&u^{-1}\big{(}6u^{-4}\langle U,A\rangle\langle A,A\rangle-4u^{-2}\langle A,A^{\prime}\rangle\big{)}U-\\ &u^{-1}\big{(}-12u^{-4}\langle U,A\rangle^{2}+4u^{-2}\langle U,A\rangle+3u^{-2}\langle A,A\rangle\big{)}A-\\ &4u^{-1}\langle U,A\rangle A^{\prime}+u^{-1}A^{\prime\prime}.\end{split}$ (C.45)
Note that
$\alpha_{1}$ $=-6u^{-4}\langle U,A\rangle^{2}+2u^{-2}\langle U,A^{\prime}\rangle+3u^{-2}\langle A,A\rangle,$
$\Delta_{4}$ $=9u^{-8}\langle U,A\rangle^{4}-6u^{-6}\langle U,A\rangle^{2}\langle U,A^{\prime}\rangle-9u^{-6}\langle U,A\rangle^{2}\langle A,A\rangle$
$+6u^{-4}\langle U,A\rangle\langle A,A^{\prime}\rangle+u^{-4}\langle U,A^{\prime}\rangle^{2}-u^{-2}\langle A^{\prime},A^{\prime}\rangle.$
The equation (3.15) then gives that the following expression vanishes
$-24u^{-8}\langle U,A\rangle^{3}U+16u^{-6}\langle U,A\rangle\langle U,A^{\prime}\rangle U+12u^{-6}\langle U,A\rangle\langle A,A\rangle U+$
$12u^{-6}\langle U,A\rangle^{2}A-2u^{-4}\langle U,A^{\prime\prime}\rangle U-4u^{-4}\langle A,A^{\prime}\rangle U-$
$4u^{-4}\langle U,A^{\prime}\rangle A-3u^{-4}\langle A,A\rangle A-4u^{-4}\langle U,A\rangle A^{\prime}+u^{-2}A^{\prime\prime}.$
It turns out that involving the restriction $\alpha^{\prime}_{1}=0$ on this expression results in the minus $u^{-1}$-multiple of (LABEL:eq). This agrees with the fact, that the tractor equation lives in the space $TM[-1]$ while the flat Mercator equation lives in the space $TM[-2]$. Let us note that this works only in the case of a conformal sphere. In the general curved case, the Mercator equation differs from the equation coming from tractor calculations by some curvature term.